# Position: AI Competitions Provide the Gold Standard for Empirical Rigor in GenAI Evaluation

### v1

@@ 5.1 Unreleased Holdout Sets
Additionally, as holdout sets and evaluation runs must necessarily be kept private, results are not reproducible by researchers. To mitigate this, some benchmarks take a hybrid approach. For example, the FACTS Grounding Leaderboard (Jacovi et al., 2025) publishes half the test set publicly which enables partial reproducibility and better understanding of the benchmark. A model’s performance can be compared on private and public parts of test sets to identify models that may have (intentionally or not) trained on test.
@@ 5.2 Dynamic Benchmarks
LiveBench (White et al., 2024), LiveCodeBench (Jain et al., 2024), and SWE-Bench (Jimenez et al., 2024) benchmarks frequently update test sets from sources that refresh naturally over time. For example, LiveBench test set questions are updated weekly from sources such as newly published news articles or papers on arXiv. By testing models only with very recent data, benchmarks can mitigate the risk that such data was included in model training. LiveBench also does not release the most recently added test data so that a significant percentage of questions are always heldout.
@@ 5.3 Community Benchmarks
Community benchmarks are difficult to build and maintain. To evaluate many models, the number of votes required can be very large, necessitating a large and constant pool of voters. Community benchmarks also don’t work for all tasks, for example, tasks that require specialized knowledge or which might take humans many days to complete will not scale to human rating. Community benchmarks are also necessarily biased by any sampling effects; the diversity and distribution of voters can affect results and great attention and care is required to filter out low quality, duplicate, or contaminated results.
@@ 6 AI Competitions as Structural Solution
AI Competitions typified by platforms like Kaggle and others offer an “embarrassingly parallel” structure to empirical evaluation shown in Figure 2 that hearkens back to the classic MapReduce structure from parallel computing (Dean & Ghemawat, 2008). In this structure, independent teams of researchers—often numbering in the thousands—each compete to solve a given problem, and in so doing create an evaluation of many different approaches in one massive parallel effort. Here we show ways that this structure offers useful benefits to the the problem of GenAI evaluation writ large.
@@ 6.1 Parallelization Improves Robustness
In addition, competition platforms such as Kaggle can serve as trusted keepers of hidden test data by running isolated code competitions, where competitors submit their models to be run on an isolated, secure backend without access to internet. By evaluating all models securely offline, competitions platforms can guarantee no hidden test data is leaked.
Finally, competitions hosted on large community platforms offer additional non-structural characteristics which represent best practices the industry should adopt to further improve empirical rigor. Competitions encourage or often require open sharing of code, data, and experimental details, including both successes and failures. This transparency facilitates reproduction of results, fosters trust in new baselines, and accelerates the dissemination of knowledge within the research and practitioner communities.
@@ 6.2 Leak-Proof Competition Structures
We provide some examples of competitions that demonstrate the feasibility of leak-proof evaluation design. Employing strategies such as prospective ground truth, novel task generation, and post-deadline data collection, generally combined with test data that is directly inaccessible to competitors, competitions can provide a robust and reliable platform for novel evaluation of GenAI models. These best practices should be considered and adapted as blueprints for future competition and benchmark design.
@@
The Critical Assessment of protein Function Annotation (CAFA) 5 challenge (Friedberg et al., 2023) is an example of a competition that uses a prospective ground truth to mitigate leakage. The competition took as its test set proteins whose sequences were known, but whose functional annotations had not yet been determined in a wet lab. Nearly two thousand participants across 1,625 independent teams therefore effectively developed models predicting the function of a set of proteins without any ground truth yet available to any human or model during an active training phase. Months later, the final evaluation was determined following a ”curation phase” on the basis of newly published protein functions. This novelty makes the competition reasonably leak-proof.
@@ 7 Recommendations for the Field
Adopt and improve on the anti-cheating structures from AI Competitions to improve standard practice for GenAI evaluations. Furthermore, as a field, we can learn from the best practices that have been developed by AI Competitions – the techniques and practices that have been created to combat intentional cheating by bad actors are equally valuable in creating evaluation structures that combat unintentional issues such as leakage and contamination that may invalidate empirical results. A cheat-proof structure is one that provides assurance to researchers that they will not accidentally cheat themselves. We also need to augment and further improve these structures, for example by creating a field-wide standard that major API-based models agree to follow to explicitly avoid collecting or training on data that may appear in evaluations.
@@ 8 Alternative Views
All position papers should consider opposing views, and ours is no exception. One reasonable alternative view is that the current state of benchmarking is proceeding well without the need for additional intervention. The many new static benchmarks are appearing on platforms like Hugging Face, OpenML, and Kaggle on a near-daily basis and may serve as the steady stream of novel tasks that we described as necessary for the field. While we applaud all efforts to create new benchmarks, we do fundamentally believe that static benchmarks should be considered to have been effectively invalidated once they have been published, and thus it is the time component of AI Competitions that provides unique additional value.

### v2

@@ 5.1 Unreleased Holdout Sets
Additionally, as holdout sets and evaluation runs must necessarily be kept private, results are not reproducible by researchers. To mitigate this, some benchmarks take a hybrid approach. For example, the FACTS Grounding Leaderboard (Jacovi et al., 2025) publishes half the test set publicly which enables partial reproducibility and better understanding of the benchmark. A model’s performance can be compared between the private and public parts of test sets to identify models that may have (intentionally or not) trained on test data or leakage.
@@ 5.2 Dynamic Benchmarks
LiveBench (White et al., 2025), LiveCodeBench (Jain et al., 2024), and SWE-Bench (Jimenez et al., 2024) benchmarks frequently update test sets from sources that refresh naturally over time. For example, LiveBench test set questions are updated weekly from sources such as fresh news articles or papers on arXiv. By using only very recent data, benchmarks can mitigate if not eliminate the risk that such data was included in model training. LiveBench also does not release the most recently added test data so that a significant percentage of questions are always private and heldout.
@@ 5.3 Community Benchmarks
Community benchmarks are difficult to build and maintain. To evaluate many models, the number of votes required can be very large, necessitating a large and constant pool of voters. Community benchmarks also don’t work for all tasks. For example, tasks that require very specialized knowledge or which might take humans many days to verify will not scale to human rating. Community benchmarks are also necessarily biased by any sampling effects; the diversity and distribution of voters can affect results and great attention and care is required to filter out low quality, duplicate, or contaminated results.
@@ 6 AI Competitions as Structural Solution
AI Competitions typified by platforms like Kaggle and others offer an “embarrassingly parallel” structure to empirical evaluation shown in Figure 2 that hearkens back to the classic MapReduce structure from parallel computing (Dean & Ghemawat, 2008). In this structure, independent teams of researchers—often numbering in the thousands—each compete to solve a given problem, and in so doing create an evaluation of many different approaches in one massive parallel effort. Here we show ways that this structure offers useful benefits to the problem of GenAI evaluation at large.
@@ 6.1 Parallelization Improves Robustness
In addition, competition platforms such as Kaggle can serve as trusted keepers of hidden test data by running isolated code competitions, where competitors submit their models to be run on an isolated, secure backend without network access. By evaluating all models securely offline, competitions platforms can guarantee no hidden test data is leaked.
Finally, competitions hosted on large community platforms offer additional non-structural characteristics which represent best practices the industry should adopt to further improve empirical rigor. Competitions encourage or often require open sharing of code, data, and experimental details, including both successes and failures. Competitors are often more motivated by the status and recognition gained from sharing valuable and insightful resources and ideas than by winning prizes. In fact, the median number of forum messages per Kaggle featured competition last year was 1,400 (data is available in the Meta Kaggle dataset (Risdal & Bozsolik, 2022). This transparency facilitates reproduction of results, fosters trust in new baselines, and accelerates the dissemination of knowledge within the research and practitioner communities.
@@ 6.2 Leak-Proof Competition Structures
We provide some examples of competitions that demonstrate the feasibility of leak-proof evaluation design. Employing strategies such as prospective ground truth, novel task generation, and post-deadline data collection, generally combined with test data that is directly inaccessible to competitors, competitions can provide a robust and reliable platform for novel evaluation of GenAI models. Overall, the main method for creating leak-proof competitions involves evaluating models based on data that does not exist at training time. These best practices should be considered and adapted as blueprints for future competition and benchmark design.
@@
The Critical Assessment of protein Function Annotation (CAFA) 5 challenge (Friedberg et al., 2023) is an example of a competition that uses a prospective ground truth to mitigate leakage. The competition took as its test set proteins whose sequences were known, but whose functional annotations had not yet been determined in a wet lab. Nearly two thousand participants across 1,625 independent teams therefore effectively developed models predicting the function of a set of proteins without any ground truth yet available to any human or model during an active training phase. Months later, the final evaluation was determined following a “curation phase” on the basis of newly published protein functions. This novelty makes the competition reasonably leak-proof.
@@ 7 Recommendations for the Field
Areas for fruitful meta-analysis include but are not limited to:
  • Studying Problems: Exploring the nature and number of problems addressed via AI Competitions. Are there underrepresented problems? How well do AI Competitions connect to real-world performance on similar problems (i.e., ecological validity)?
  • Summarizing Results: Aggregating and synthesizing trends across AI Competition results within similar domains or against similar tasks, including investigating trends (including longitudinal) in techniques, tools, models, etc. One example of this is the annual State of ML Competitions report (Carlens, 2025).
  • Improving AI Competition Design: Meta-analyses of AI competition design including choice of objective evaluation metrics, anti-cheating and anti-contamination mechanisms, etc. to help the field develop and adopt better methodologies for AI evaluation and model building.
  • Analyses of Team and Social Collaborative-Competitive Dynamics: Identifying patterns from the “parallelized attempts”, i.e., the teams working together and against one another in competition. What characterizes successful teams and individuals across competitions?
Adopt and improve on the anti-cheating structures from AI Competitions to improve standard practice for GenAI evaluations. Furthermore, as a field, we can learn from the best practices that have been developed by AI Competitions. The techniques and practices that have been created to combat intentional cheating by bad actors are equally valuable in creating evaluation structures that combat unintentional issues such as leakage and contamination that may invalidate empirical results. A cheat-proof structure is one that provides assurance to researchers that they will not accidentally cheat themselves. We also need to augment and further improve these structures, for example by creating a field-wide standard that major API-based model creators agree to follow to explicitly avoid collecting or training on data that may appear in evaluations.
@@ 8 Alternative Views
All position papers should consider opposing views, and ours is no exception. One reasonable alternative view is that the current state of benchmarking is proceeding well without the need for additional intervention. The many new static benchmarks appearing on platforms like Hugging Face, OpenML, and Kaggle on a near-daily basis may serve as the steady stream of novel tasks that we described as necessary for the field. While we applaud all efforts to create new benchmarks, we do fundamentally believe that static benchmarks should be considered to have been effectively invalidated once they have been published, and thus it is the time component of AI Competitions that provides unique additional value.
Another possible critique of AI Competitions compared to “evergreen” static benchmarks is that an artificial deadline may prohibit valuable submissions. We’ve found that every time we’ve ensembled submissions, we obtain little-to-no improvement to top ranked solutions. In other words, competitions at least on Kaggle extract (near) maximal signal from the data within the competition’s constraints.
Furthermore, AI Competition hosts are strongly incentivized to design good evaluation metrics, and we observe that outcomes where solutions correlate with real-world performance are more likely. For example, in the OpenVaccine Challenge (Wayment-Steele et al., 2022), competitors improved the state-of-the-art in mRNA vaccinate degradation rate prediction by 25% within just 4 weeks, and the hosts further validated that the solutions generalized to much longer RNA sequences not seen as part of the competition dataset.
@@
Acknowledgments
The authors would like to express deep gratitude to those who have played an instrumental role in building and evolving Kaggle as an AI Competitions platform that sets the gold standard today for GenAI evaluation: Anthony Goldbloom, Jeremy Howard, Ben Hamner, Jeff Moser, and many other members of the Kaggle team, past and present. We’re equally grateful to the many hundreds of hosts and partners who have brought diverse and challenging problems to the AI community through Kaggle. Their contributions and experiences have led to many insights and learnings shared in this paper.
We’ve also benefited from discussion with members of Kaggle’s Research Advisory Board: Isabelle Guyon, Jonathan Frankle, Sara Hooker, and Joelle Pineau.
Kaggle is a wholly owned subsidiary of Google, Inc. and we gratefully acknowledge Google and Alphabet for their ongoing support for this work.
Finally, we are thankful to the AI community who has participated in Kaggle Competitions and shared their knowledge openly with the world.